Sir please tell me that
X-2/3=0
Can we write it as
X=+3/2

Answer by Alan3354(69443)

Sir please tell me that
X-2/3=0
Can we write it as
X=+3/2
------------------
No.
If you mean
x - (2/3) = 0
Add 2/3
x = 2/3
================
If you mean
(x-2)/3 = 0
x-2 = 0
x = 2